OpenStudy (anonymous):

Solve the system. y = x + 2 and y = -x + 0 (1, -1) (-1, 1) (1, 3) (-1, 3) no solution

myininaya (myininaya):

We want to find where the lines intersect so find where both y's are the same by setting x+2 equal to -x+0 So solve x+2=-x for x and let me know what you get

OpenStudy (anonymous):

-1 ?

myininaya (myininaya):

Right. x=-1 Now replace x with -1 in one of the equations you began with to find y

myininaya (myininaya):

You can use either y=x+2 or y=-x

OpenStudy (anonymous):

So would the answer be no soloution?

myininaya (myininaya):

? Did you plug in -1 for x into one of the equations you started with?

OpenStudy (anonymous):

Ohh, so would it be (1,-1)

myininaya (myininaya):

You said x was -1

myininaya (myininaya):

What is y?

myininaya (myininaya):

You have y=x+2 or y=-x that you can use to find y

myininaya (myininaya):

Just choose one of the equations y=-x

myininaya (myininaya):

We have x=-1 Replace x with -1

myininaya (myininaya):

What do you get for y?

myininaya (myininaya):

y=-(-1)=?

myininaya (myininaya):

So after you figure out what y is which is -(-1) <---simplify this of course Then (x,y) is the your ordered pair

OpenStudy (theviper):

\[\Huge{\color{red}{\star \star}{(-1,1)}}\]

OpenStudy (theviper):

First find the value of y in any eq than put this value in other eq.:)